our class planned a holiday party for disadvantaged kids. Some of us baked cookies for the party. On the day of the party,we found we could divide the cookies into packets of two,three,four,five,or,six and have just one cookie left over in each case. If we divided them into packets of seven, there would be no cookies left over. What is the the least number of cookies the class could have baked?

Answer by richwmiller(17219)

721/2=360+1
721/3=240+1
721/4=180+1
721/5=144+1
721/6=120+1
721/7=103
I have 721 as the least but I think there should a smaller number.